Students at a high school know that January means preparation for midterms. About 11 out of 20 students usually increase their amount of studying in preparations for midterms. What is the probability that less than 40% of a 45 person sample actually increases its studying habits?
Thank you for your help! I have a whole sheet with similar problems that I don't understand, but I'm sure that by understanding this one, I can solve the rest, so thanks again! :)

Answer by stanbon(57347)

Students at a high school know that January means preparation for midterms. About 11 out of 20 students usually increase their amount of studying in preparations for midterms. What is the probability that less than 40% of a 45 person sample actually increases its studying habits?
----
mean = 11/20 = 0.55
----
z(0.40) = (0.40-0.55)/sqrt[0.55*0.45/45] = -2.0226
---
P(p < 0.4) = P(z < -2.0226) = 0.0216
=======================================================
Cheers,
Stan H.
=================